OpenStudy (anonymous):

(7k-1)^2=-125 find the imaginary number solutions

OpenStudy (anonymous):

\[\left\{k=\frac{1}{7} \left(1-5 i \sqrt{5}\right),k=\frac{1}{7} \left(1+5 i \sqrt{5}\right)\right\} \]